OpenStudy (anonymous):

Factor completely: 6x2 + 19x + 3 A. (6x + 1)(x + 3) B. (3x + 1)(2x + 3) C. (3x + 3)(2x + 1) D. (6x + 3)(x + 1)

OpenStudy (anonymous):

6*3=18 18+1=19 18*1=18 write 19x= 18x+1x and make factors.

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

Thanks Alex

OpenStudy (anonymous):

A